### v2

# Dynamic Multi-Expert Aggregation

Abstract
Combining pre-trained expert models offers substantial potential for scalable multimodal reasoning, but building a unified framework remains challenging due to the increasing diversity of input modalities and task complexity. For instance, medical diagnosis requires precise reasoning over structured clinical tables, while financial forecasting depends on interpreting plot-based data to make informed predictions. To tackle this challenge, we introduce MEXA, a training-free framework that performs modality- and task-aware aggregation of multiple expert models to enable effective multimodal reasoning across diverse and distinct domains. MEXA dynamically selects expert models based on the input modality and the task-specific reasoning demands (i.e., skills). Each expert model, specialized in a modality-task pair, generates interpretable textual reasoning outputs. MEXA then aggregates and reasons over these outputs using a Large Reasoning Model (LRM) to produce the final answer. This modular design allows flexible and transparent multimodal reasoning across diverse domains without additional training overhead. We extensively evaluate our approach on diverse multimodal benchmarks, including Video Reasoning, Audio Reasoning, 3D Understanding, and Medical QA. MEXA consistently delivers performance improvements over strong multimodal baselines, highlighting the effectiveness and broad applicability of our expert-driven selection and aggregation in diverse multimodal reasoning tasks.
MEXA: Towards General Multimodal Reasoning with
Dynamic Multi-Expert Aggregation
Shoubin Yu1,thanks: Equal contribution.   Yue Zhang1,∗   Ziyang Wang1 Jaehong Yoon1,2Mohit Bansal1 1UNC Chapel Hill   2Nanyang Technological University https://github.com/Yui010206/MEXA
1 Introduction
The rapid advancement of multimodal learning has significantly improved AI systems’ ability to understand, reason, and interact with the real world, benefiting diverse tasks such as visual question answering Antol et al. (2015); Hudson and Manning (2019); Marino et al. (2019); Mathew et al. (2021); Tanaka et al. (2023); Winata et al. (2024); Wang et al. (2025); Zhang et al. (2024b); Guo et al. (2025), medical image diagnosis Lau et al. (2018); He et al. (2020); Liu et al. (2021); Azam et al. (2022); Cai et al. (2023); Bai et al. (2024), and embodied AI Brohan et al. (2023); Driess et al. (2023); Liu et al. (2024b); Durante et al. (2024); Majumdar et al. (2024); Zhang et al. (2024c). As AI systems become more integrated into complex, real-world applications, the ability to flexibly understand and reason upon heterogeneous multimodal inputs is increasingly essential. For example, a medical diagnostic system may need to interpret CT scans, extract structured information from clinical notes, and reason about temporal patterns in patient histories, while a financial forecasting system must effectively analyze and interpret plot data to make informed predictions about market trends and economic risks.
However, despite recent advances in multimodal learning, the increasing diversity and complexity of multimodal data pose significant challenges for developing a flexible and unified framework that can reason effectively across modalities and generalize across tasks requiring diverse skills. Existing multimodal architectures typically require training separate encoders tailored to each modality, along with designing complex cross-modal alignment mechanisms Tan and Bansal (2019); Li et al. (2022, 2023b); Liu et al. (2023a); Hong et al. (2023); Yu et al. (2024). While effective, these models require fine-tuning for each specific task, leading to substantial training overhead and limiting their adaptability to new modalities or tasks. Moreover, previous approaches Jaegle et al. (2021); Team (2024) often implicitly fuse multimodal inputs at early stages, restricting transparency and interpretability of the reasoning process. Instead, an ideal multimodal framework should seamlessly process inputs from any modality for any given task, dynamically directing these inputs to the most suitable multimodal expert models. This necessitates a modular design where each expert model is associated explicitly with skills corresponding to particular modality-task, thereby ensuring precise skill matching and enhancing interpretability.
In this paper, we propose Multimodal Expert Aggregator (MEXA), a novel training-free multi-expert aggregation framework that dynamically coordinates a pool of specialized experts. MEXA selectively activates and aggregates reasoning outputs from the most relevant expert models for each input instance, guided by both the input modalities (e.g., image, audio, 3D, medical scan) and the required level of reasoning demands (i.e., skills) for that instance. MEXA spans a wide spectrum of tasks, from low-level perceptual reasoning (e.g., object recognition or OCR) to high-level cognitive inference (e.g., temporal event understanding in video or diagnostic interpretation in medical imaging). Unlike existing methods that rely on monolithic, end-to-end fine-tuned multimodal models or fixed, statically defined expert pipelines, MEXA’s modular expert coordination enables flexible, interpretable, and scalable multimodal reasoning.
Specifically, we first assemble a diverse pool of expert models, each designed to handle distinct aspects of modalities and tasks, to address mainstream multimodal challenges effectively. Each expert specializes in extracting information unique to its respective skill and encoding it into a unified textual representation. Then, to effectively coordinate these experts, we design an expert selection module that employs a Multimodal Large Language Model (MLLM) as a versatile router within our framework. This router dynamically selects and activates the appropriate experts by analyzing the modality of the input data and identifying the specific skills required based on the input query and task requirements. Finally, we introduce an aggregator that reasons over the outputs of the selected experts using a Large Reasoning Model (LRM), which excels at long-context understanding and producing complex long CoT reasoning during LRM inference. Instead of relying on heuristic merging, the aggregator systematically integrates the complementary information from each expert and infers the final answer based on the full context of the target task.
Figure 1: Overview of the MEXA Architecture. Given the input task context and question, MEXA first employs an MLLM router (Sec.˜3.2.2) to select the appropriate experts based on input modality and required reasoning skills. The aggregator (Sec.˜3.2.3) then reasons over the outputs from the selected experts to generate the final answer.
We perform extensive evaluations of our framework on a diverse set of challenging multimodal benchmarks containing various modalities and requiring diverse skills, including expert knowledge heavy video reasoning (Video-MMMU Hu et al. (2025a)), audio QA (MMAU Sakshi et al. (2024)), 3D scene understanding (SQA3D Ma et al. (2023a)), and medical imaging-based QA (M3D Bai et al. (2024)). Experimental results show that our proposed method consistently outperforms strong multimodal baseline models across all evaluated benchmarks, with accuracy gains of +5.7% on Video-MMMU, +12.2% on MMAU, +1.7% on SQA3D, and +1.6% on M3D in the corresponding metrics, demonstrating the effectiveness, flexibility, and robustness of our expert-driven multimodal aggregation framework. We further conduct ablation studies on expert-selection router design and LLM aggregator to provide more insights for future work. Our contributions are summarized as follows:
  • We present MEXA, a training-free, modality-extensible, and flexible framework that handles general multimodal reasoning via dynamic expert selection and aggregation.
  • MEXA achieves strong performance on challenging multimodal benchmarks across diverse modalities, against specialized models as a more general framework.
  • We conduct ablation studies and analyses based on our MEXA framework, providing insights into the design choices and effectiveness of each component.
2 Related Work
Mixture of Multiple Expert Models.
Recently, the Mixture-of-Experts (MoE) paradigm Shazeer et al. (2017); Zoph et al. (2022); Chen et al. (2023b); Zhou et al. (2022); Dai et al. (2022); Zhang et al. (2023); Chowdhury et al. (2023); Lee et al. (2024), which integrates multiple specialized parametric expert modules, has been widely adopted across a range of machine learning tasks to leverage complementary capabilities. These methods typically employ sparsely-gated mechanisms that activate only a subset of expert modules during each forward pass, improving computational efficiency and scalability. Building on this motivation, recent advances have explored leveraging a mixture of expert models (or agents(Li et al., 2024b; Wang et al., 2024; Chen et al., 2025; Li et al., 2025; Cao et al., 2025) that combine independently pre-trained models across diverse knowledge domains, moving beyond the traditional layer-wise sparse activation used in MoE-based approaches. Rather than operating as interchangeable sub-modules within a single model, such agent-based (or model-based) mixtures selectively aggregate the outputs of independently trained models through dynamic routing mechanisms to solve tasks. This design allows for more flexible and targeted knowledge utilization, promoting improved generalization across tasks and greater extensibility. However, existing approaches primarily focus on single-modality settings Chen et al. (2025) or lack support for complex reasoning tasks Li et al. (2024b); Wang et al. (2024); Li et al. (2025); Cao et al. (2025); Zhang et al. (2025). In contrast, our method extends this line of work by scaling expert modularity to handle heterogeneous multimodal inputs and reasoning, enabling more versatile and adaptive problem-solving across domains.
Many-Modal Understanding and Reasoning.
Real-world environments are increasingly dynamic, requiring the need for AI systems to perceive and process a broader range of modalities beyond unimodal learning. This growing demand has led to advances in models that integrate diverse signals such as images, audio, text, and 3D point clouds to improve learning and generalization. Among these, Vision-Language Models (VLMs) (Huang et al., 2023; Li et al., 2023a; Gong et al., 2023; Chen et al., 2023a) support tasks like speech recognition and audio captioning by fusing acoustic and textual information. Similarly, 2D-3D Joint Models (Li et al., 2020; Hou et al., 2021, 2023; Lei et al., 2024) combine spatial and geometric features for enhanced 3D scene understanding. However, these models are typically limited to fixed modality-task combinations and struggle to generalize or scale to novel modality inputs. This motivates the development of flexible many-modality systems (Zellers et al., 2022; Han et al., 2023; Li et al., 2023b; Girdhar et al., 2023; Liu et al., 2023b; Yu et al., 2024) that adapt to heterogeneous inputs and reasoning demands across domains. Yet, most existing systems are primarily designed to fuse multiple input modalities and learn perception and reasoning implicitly within a unified architecture. While effective for many tasks, these approaches often struggle with scalability and tend to overlook the explicit use of LLMs’ advanced reasoning capabilities. Our modular framework selects and coordinates expert models through an LLM-driven router and aggregator, directly leveraging reasoning abilities that are insufficiently exploited by unified multimodal architectures. Another line of work, programming-based methods Lu et al. (2023); Surís et al. (2023), relies on pre-defined functions or in-context examples (few-shot prompts) and executes reasoning in a code format. In contrast, our MEXA does not require in-context examples and operates entirely in a training-free, zero-shot setting. MEXA performs deep reasoning using a Large Reasoning Model rather than code execution, allowing for free-form input rather than rigid programmatic steps. This flexibility makes MEXA more extensible and generalizable to a wider range of tasks and modalities, including audio, 3D, and medical tasks, which are beyond the scope of ViperGPT Surís et al. (2023) and Chameleon Lu et al. (2023).
3 MEXA: Dynamic Multi-Modal Expert Aggregation
We first highlight the challenges faced by general-purpose multimodal reasoning systems in Sec.˜3.1 and introduce our skill-specialized expert aggregation framework as a solution in Sec.˜3.2. Within Sec.˜3.2, we first describe the organization of a pool of expert modules and our strategy for extracting expertise-specific text representations from the specialized experts(Sec.˜3.2.1). Next, we introduce an expert selection module that dynamically determines which experts to activate based on the input query, target modality, required skill, and reasoning complexity (Sec.˜3.2.2). Finally, we present our expert aggregation strategy, which integrates the outputs from the selected expert to generate a final answer (Sec.˜3.2.3).
3.1 General-Purpose Multimodal Reasoning
Existing general-purpose multimodal reasoning systems Han et al. (2023); Liu et al. (2023b); Yu et al. (2024) typically rely on fixed fusion architectures that process raw multimodal inputs within a shared representation space. These systems aim to implicitly learn both multimodal alignment and reasoning capabilities within a single end-to-end model. Formally, such a system can be framed as a question-answering model that takes a natural language question $Q_{\text{ques}}$ and a collection of supported modalities, such as 2D images, 3D point clouds, videos, text, and audio, denoted as $\mathcal{M}=\{m_{1},m_{2},\dots,m_{n}\}$, where $n$ is the number of input modalities as input. The input modalities are first encoded and fused into a joint representation space, which is then processed by a monolithic reasoning model trained to reason over this fused representation. The model produces a contextually grounded answer $A$, defined as:
$A=f_{\text{fuse}}(f_{\text{enc}}(\mathcal{M}),Q_{\text{ques}}),$ (1)
[…]
3.2 Skill-Specialized Mixture of Expert Models
To overcome the limitations of existing general-purpose multimodal models in handling diverse modalities and tasks, we propose MEXA, a flexible and expert-driven multimodal reasoning framework. Instead of processing all modality inputs uniformly through fixed architectures, MEXA dynamically routes queries to a set of specialized expert models, selected based on input modality and task complexity. Formally, our system takes a question $Q_{\text{ques}}$ as input and generates an answer $A$ in response. To accomplish this, we first employ an expert selection module that serves as a router (denoted as Router), which selects a subset of expert models, denoted as $E_{s}$. The outputs from the selected experts are then passed to a reasoning module, functioning as an aggregator (denoted as Aggregator), which integrates and reasons over the information extracted by expert models to generate the final answer $A$:
$\begin{split}A=\texttt{Aggregator}(\{E_{s}\mid E_{i}\in\texttt{Router}(\\ Q_{\text{ques}},\cdots)\}).\end{split}$ (2)
In contrast to Equation 1, which implicitly fuses all modalities within a single architecture, our method enables specialization and modular design. It supports dynamic adaptation by selecting relevant experts based on the task, where each expert plays a well-defined role.
3.2.1 Design Principles for the Expert Pool
Our expert pool is designed based on the following key principles.
Design Principle 1: Task-Aware and Modality-Sensitive Reasoning.
To ensure that our expert pool is both modular and broadly applicable, we construct it by analyzing the modalities and skills commonly required across diverse multimodal tasks. While our approach is designed to be task-agnostic and extensible, we use the Video-MMMU dataset Hu et al. (2025a) as a representative case study to guide the initial design and evaluation. This benchmark includes videos from a wide range of domains such as medicine, mathematics, and the arts, allowing us to capture a diverse set of modalities and task combinations. For each domain, we systematically identify the core perception and reasoning skills required to answer the associated questions, ensuring that our expert pool is both modular and generalizable across tasks. This design allows our framework to scale beyond any single dataset and adapt to new domains with minimal effort.
In Fig. 1, we illustrate the complete set of expert modules assembled in our framework. Specifically, we categorize these experts into four distinct types: (1) Perceptual Experts, specialized in extracting visual information directly from images, videos, audio, 3D point clouds and medical images; (2) Text Extraction Experts, which leverage optical character recognition (OCR) to distill textual content from visuals such as slides or embedded text; (3) Structured Experts, designed to analyze structured visual data such as charts, tables, and diagrams; and (4) Mathematical Reasoning Expert, focused on interpreting and solving questions related to mathematical and geometric equations presented in LaTeX format. Furthermore, for each expert category, we design skill-specialized experts that target distinct reasoning demands across modalities and task types. We focus on aligning each expert’s functionality with the specific requirements of the task. For instance, for 2D image perception, we design experts aligned with skills corresponding to different levels of reasoning granularity: a fine-grained image description expert, responsible for generating comprehensive captions covering multiple visual elements and their relationships, and a concise summarization expert, which produces brief, high-level summaries highlighting only the salient visual content. Within the domain of 3D scene understanding, we develop two experts with specialized skills for different task demands: a general 3D scene expert, which provides descriptions capturing the overall spatial layout and major objects in the scene, and a situated 3D expert, designed to generate detailed, viewpoint-grounded descriptions reflecting the specific perspective of an embodied agent positioned within the environment.
Design Principle 2: Unified Textual Representation.
We design expert models that convert diverse modality-specific inputs into a shared textual representation, facilitating the integration of heterogeneous multimodal data. To achieve this, we leverage a series of state-of-the-art captioning models combined with modality- and skill-specific prompting strategies. Each expert in our framework functions as a captioner (descriptor), converting modality- and skill-specific information into a natural language format. This textual abstraction further enables interpretable perception and supports reasoning at varying levels of complexity. Formally, given a set of supported input modalities $\mathcal{M}$, the objective is to obtain a unified textual representation $x$ by converting the relevant modality-specific input into natural language. This is accomplished through a set of modality-specific expert functions $E_{i}$, and textual output from each expert is denoted as:
$x_{i}=E_{i}(m_{i},p_{i}),$ (3)
where $p_{i}$ is a carefully designed prompt that reflects the reasoning objective for different modalities and reasoning demands.
Figure 2: Expert distributions selected by MEXA across different benchmarks, covering video (Video-MMMU), audio (MMAU), 3D (SQA3D), and medical imaging (M3D).
3.2.2 Expert Selection Module
After constructing the pool of expert models, the next step is to determine which experts to activate for a given task. We design an expert selection module that leverages a multimodal LLM (MLLM) as a router for different experts. This expert selection process leverages the commonsense reasoning capabilities of the MLLM, allowing it to deeply understand the semantics of the task, the relationships among modalities, and the contextual intent behind the question when deciding which experts to activate. As illustrated in Fig. 1, the router takes task context and question as input. Task context provides high-level guidance on the task type, and helps the router infer the set of skills required to answer the question and explicitly constrains the router by ensuring that only relevant experts are considered. Based on these inputs, the expert selection module activates a subset of expert models most suited to addressing the question.
Formally, given question $Q_{\text{ques}}$, a task description $T_{r}$, also a set of available modality inputs $\mathcal{M}$, the MLLM router selects a subset of expert models based on needed modality and tasks. We denote the selected experts models as $\{E_{s}\}_{s\in\mathcal{S}}$, where $\mathcal{S}\subseteq\mathcal{M}$. Once the experts are selected, each expert $\{E_{s}\}$ receives the specific prompt $p_{s}$ and transforms the input into a natural language description $x_{s}$. The resulting set of expert-generated textual output is represented as :
[…]
where $\mathcal{T}$ represents the set of intermediate textual outputs generated by the selected experts, serving as a unified and interpretable representation that is later aggregated to produce the final answer.
3.2.3 Aggregating Expert Information via Reasoning
Once the selected experts have generated modality- and task-specific textual representations, we employ a reasoning module based on LRM, selected for its strong long-context reasoning and deep inference capabilities. This LRM-based aggregator systematically integrates the diverse outputs from the selected experts and reasons over them to generate the final answer. Formally, the expert outputs are represented as $\mathcal{T}=\{t_{1},t_{2},\dots,t_{k}\}$, where $k$ denotes the number of selected expert. The aggregator inputs the entire set $\mathcal{T}$, along with the task description for the expert, which is denoted as $T_{a}$, and generates the final answer, denoted as:
$\begin{split}A=\texttt{Aggregator}(\mathcal{T};T_{a})\end{split}$ (5)
By aggregating expert-generated textual outputs and performing reasoning over these unified representations, MEXA enables more accurate, interpretable, and task-aligned answers, demonstrating clear advantages over single end-to-end models that process implicit representations without explicit decomposition or modular coordination.
4 Experimental Results
4.1 Experimental Setup
We use GPT-4o Hurst et al. (2024) as our multimodal expert selection module to dynamically select the most relevant experts based on the input modalities and task requirements. Subsequently, we employ Deepseek as our aggregator to reason over the expert-generated textual information to obtain the final answer.
Method Overall Results by Track Results by Discipline
Perception Comprehension Adaptation  Art. Biz. Sci. Med. Hum. Eng.
Random Choice $14.0$ $12.0$ $14.0$ $16.0$  $11.1$ $12.9$ $12.1$ $22.5$ $10.5$ $13.6$
Open-source LMMs
LLaVA-OneVision-72B Li et al. (2024a) $48.3$ $59.7$ $42.3$ $43.0$  $61.9$ $46.21$ $40.2$ $54.3$ $60.0$ $44.0$
LLaVA-Video-72B Zhang et al. (2024a) $49.7$ $59.7$ $46.0$ $43.3$  $69.8$ $44.7$ $41.7$ $58.9$ $57.1$ $45.1$
Aria Li et al. (2024c) (8 $\times$ 3.5B) $50.8$ $65.7$ $46.7$ $40.0$  71.4{71.4} $47.7$ $44.7$ $58.9$ $62.9$ $43.7$
Proprietary LMMs
Gemini 1.5 Flash Team et al. (2024) $49.8$ $57.3$ $49.0$ $43.0$  $63.5$ $53.0$ $43.2$ $49.6$ $59.1$ $45.7$
Gemini 1.5 Pro Team et al. (2024) $53.9$ $59.0$ $53.3$ $49.3$  $57.1$ $59.1$ $49.1$ $57.4$ $58.1$ $50.3$
GPT-4o Hurst et al. (2024) $61.2$ $66.0$ $62.0$ 55.7{55.7}  $69.5$ $66.9$ $51.6$ 64.8{64.8} $69.5$ $57.1$
Claude-3.5-Sonnet Anthropic (2024) $65.8$ 72.0{72.0} 69.7{69.7} 55.7{55.7}  $66.7$ 75.0{75.0} 56.1{56.1} $58.1$ 75.2{75.2} 66.1{66.1}
Human Expert $74.4$ $84.3$ $78.7$ $60.3$  $81.0$ $78.8$ $74.2$ $70.5$ $84.8$ $69.9$
MEXA (Ours) $\mathbf{71.5}$ $\mathbf{77.0}$ $\mathbf{76.7}$ $\mathbf{60.0}$  $\mathbf{76.2}$ $\mathbf{77.0}$ $\mathbf{63.8}$ $\mathbf{75.8}$ $\mathbf{78.1}$ $\mathbf{67.7}$
Table 1: Video-MMMU Evaluation Results across three cognitive tracks (Perception, Comprehension, Adaptation) and six disciplines (Art, Business, Science, Medicine, Humanities, Engineering). We highlight the best model performance for each metric.
To effectively extract textual information from diverse expert modules, we adopt the widely used captioners to generate high-quality textual descriptions across modalities. Specifically, for perceptual experts, we employ different captioning approaches to each modality. For 2D image inputs, we utilize OmniCaptioner-Qwen2-5-7B Lu et al. (2025) to generate both detailed and concise descriptions. For video modality experts, we adopt NVILA-8B Liu et al. (2024c) to generate comprehensive video-level captions. For experts specializing in 3D scene understanding, we integrate LEO-Vicuna7B Huang et al. (2024) to obtain general, top-down scene descriptions and Spartun3D-Vicuna7B Zhang et al. (2024d) for generating situated captions explicitly grounded in the agent’s viewpoint. Regarding audio modality experts, we employ Qwen-2.5-Omni Xu et al. (2025), which is adept at generating contextualized captions for both speech and music. For all non-perceptual experts, we apply OmniCaptioner-qwen-5-7B Lu et al. (2025) with prompts designed to emphasize the specific functionality of each expert. We provide detailed prompts for the selection module, aggregator, and all expert models in the Appendix.
4.2 Evaluation Datasets
We validate our framework across various challenging multimodal tasks, including Video Reasoning (Video-MMMU (Hu et al., 2025b)), Audio QA (MMAU (Sakshi et al., 2024)), 3D Situated Reasoning (SQA3D (Ma et al., 2023b)), and Medical QA (M3D (Bai et al., 2024)). We specifically chose these benchmarks because they represent diverse reasoning complexities, modality interactions, and practical application scenarios. Please see more details in the Appendix.
4.3 Quantitative Results
We evaluate MEXA on all datasets under the multiple-choice QA setting, and report performance based on standard accuracy metrics across all experiments.
Video-base Multimodal Reasoning. For the video-based multimodal reasoning capability, we evaluate MXEA on the Video-MMMU benchmark. We compare our methods with three types of baselines: open-source large multimodal models (LMMs) Li et al. (2024a, c), proprietary LMMs Hurst et al. (2024); Anthropic (2024); Team et al. (2024), and human experts Hu et al. (2025b). Results show that MEXA significantly outperforms the leading open-source LLM Li et al. (2024c) by $23.6\%$ in overall accuracy. Moreover, our method outperforms the powerful MLLM like GPT-4o Hurst et al. (2024) by $6\%$ in overall accuracy. Our method also achieves performance on par with the human expert on most metrics, showing the effectiveness of the proposed framework. Specifically, we see strong results across all six disciplines that validate the effectiveness of the combination of expert modules. For comparison, GPT-4o Hurst et al. (2024) achieves only $51.6\%$ and $57.1\%$ on science and engineering, performing $12.2\%$ and $10.5\%$ lower than MEXA, respectively. This gap highlights the advantage of explicitly aggregating and reasoning over information from multiple specialized experts, rather than relying solely on a single multimodal LLM.
Method Sound Music Speech Average
GPT-4o Hurst et al. (2024) w/ caption $39.3$ $39.5$ $58.3$ $45.7$
GAMA-7B Ghosh et al. (2024) $41.4$ $32.3$ $18.9$ $30.9$
MULLaMA-7B Liu et al. (2024a) $40.8$ $32.6$ $22.2$ $31.9$
SALAMONN-13B Tang et al. (2024) $41.0$ $34.8$ $25.5$ $33.7$
MEXA (Ours) $\mathbf{45.1}$ $\mathbf{40.7}$ $\mathbf{51.8}$ $\mathbf{45.9}$
Table 2: MMAU evaluation results across three different audio question types.
Method What Is How Can Which Others Avg.
CREMA Yu et al. (2024) $23.2$ $\mathbf{52.5}$ $34.8$ $41.5$ $34.1$ $37.5$ $37.3$
LEO Huang et al. (2024) $10.2$ $15.7$ $11.6$ $9.4$ $10.6$ $16.9$ $12.4$
Spartun3D Zhang et al. (2024d) $22.3$ $50.9$ $33.8$ $40.9$ $34.4$ $34.5$ $36.1$
MEXA (Ours) $\mathbf{23.4}$ $51.1$ $\mathbf{35.2}$ $\mathbf{42.5}$ $\mathbf{36.9}$ $\mathbf{37.9}$ $\mathbf{37.8}$
Table 3: SQA3D evaluation results across fine-grained question types.
Method Plane Phase Organ Abnormality Location Avg.
CREMA Yu et al. (2024) $14.9$ $26.7$ $15.9$ $17.3$ $13.0$ $17.2$
MiniCPM-o Yao et al. (2024) 60.7 39.2 35.2 53.0 42.0 44.7
GPT-4o 83.3 42.7 50.0 41.3 41.3 51.7
MEXA (Ours) 65.0{65.0} 48.1 60.9 $44.8$ 48.0 53.3
Table 4: M3D evaluation results across different kinds of medical QA types.
Router Aggregator Video-MMMU M3D
Qwen2.5-VL (7B) GPT-4o $57.4$ $34.7$
Qwen2.5-VL (7B) DeepSeek $70.4$ $45.8$
GPT-4o GPT-4o $58.9$ $48.2$
GPT-4o DeepSeek $\mathbf{71.5}$ $\mathbf{53.4}$
Table 5: Ablation study on MLLM router and LLM aggregator.
Audio-based Multimodal Reasoning. Tab.˜3 presents the results of the MMAU benchmark, evaluating models on three audio QA types: Sound, Music, and Speech. Our proposed MEXA consistently outperforms all audio large language models, including GAMA, MuLLaMA, and SALAMONN, across all categories. Notably, MEXA achieves a substantial improvement of $4.1\%$, $5.9\%$, and $26.3\%$ over the best-performing Audio LLM baselines, demonstrating its strong and generalizable ability to understand diverse non-visual sensory inputs, such as audio. In contrast, existing models struggle particularly with complex music and speech QA tasks. Compared to GPT-4o with generated captions using Kim et al. (2024); Liu et al. (2024a); Radford et al. (2023), MEXA achieves comparable overall performance while maintaining a more balanced accuracy across all audio types. These results validate the effectiveness of our modular expert-based approach in addressing the diverse challenges of audio understanding.
3D Scene Understanding. Results on the SQA3D benchmark shown in Tab.˜3 demonstrate the effectiveness of our method in 3D situated reasoning tasks. By integrating situated captions and general scene descriptions through our routing and aggregation system, MEXA consistently improves accuracy across all question types. Specifically, our approach achieves a $2\%$ accuracy gain over SOTA 3D-based LLMs, which rely on general-purpose encoder-fusion strategies with static unified architectures. This improvement highlights the advantage of our dynamic, modality-aware expert selection and aggregation framework in the 3D domain.
Medical QA. To demonstrate the effectiveness of our approach beyond general domains such as audio and video, we additionally evaluate MEXA on the medical video QA domain using the M3D benchmark. This benchmark includes five distinct medical QA types: Plane classification, Phase recognition, Organ identification, Abnormality detection, and Location estimation. As shown in Tab.˜5, MEXA significantly outperforms strong general-purpose MLLM, CREMA, MiniCPM-o, and GPT-4o, which directly process 3D scan images as input, and observe that MEXA consistently achieves superior performance across most categories, improving average accuracy by 36.1%, 8.7% and 1.6%, respectively. These results underscore MEXA’s robust capability in medical content understanding, even without extensive domain-specific fine-tuning.
4.4 Ablation Study and Analysis
Ablation on Router and Aggregator.
Table 5 presents ablation results for different router and aggregator configurations. We observe that, in addition to the reasoning strength of the model itself, the correct deployment and targeted usage of these models significantly contribute to performance. Specifically, we compare the effectiveness of Qwen2.5-VL and GPT-4o as multimodal expert selection modules, alongside GPT-4o and DeepSeek as aggregators, across two benchmarks: Video-MMMU and M3D. We observe that GPT-4o consistently outperforms Qwen2.5-VL when utilized as the router, indicating its superior capability in multimodal expert selection. Furthermore, DeepSeek demonstrates better performance than GPT-4o in the aggregator role, underscoring its effectiveness in reasoning over the outputs from specialized experts. This superior performance indicates that our aggregator requires strong reasoning abilities, including the effective handling of long-context information and nuanced inference across diverse expert outputs, in which DeepSeek particularly excels.
Expert Distributions.
In Fig.˜2, we visualize the expert selection distributions from MEXA’s expert selection module. For the Video-MMMU benchmark, multiple experts are frequently selected, aligning well with the benchmark’s inherent multimodal and multidisciplinary nature, which demands diverse skills for comprehensive video understanding. In the MMAU audio reasoning benchmark, the audio-specific experts (music, audio, and speech) are consistently activated, with their selections evenly distributed, reflecting balanced reliance on all audio modalities. For the 3D reasoning task (SQA3D), the situated 3D and general 3D scene experts are predominantly selected, with occasional inclusion of the image descriptor expert. Finally, for the medical dataset (M3D-VQA), the CT scan expert is primarily selected, complemented occasionally by the medical image expert. These observations confirm that our selection module effectively and adaptively matches experts to the specific modality requirements and reasoning contexts of each task. We provide qualitative examples of Video-MMMU and SQA3D in the Appendix.
5 Conclusion
In this paper, we introduce MEXA, a dynamic multimodal reasoning framework that adaptively leverages a diverse set of expert modules with the skills for specific modalities and tasks. By effectively integrating an expert selection module and aggregation mechanism, our method achieves superior performance across a variety of challenging benchmarks, including Video Reasoning, Audio QA, 3D Situated Reasoning, and Medical QA. Extensive experimental analyses and visualization results demonstrate the clear advantages of our dynamic and modality-aware design, highlighting the significance of selectively combining complementary expert skills. We believe that our framework provides a robust foundation for future research in generalized multimodal reasoning, enabling models to tackle increasingly complex real-world multimodal applications.
Limitations
Our framework achieves competitive reasoning across diverse domains by leveraging modular expert outputs. However, since it relies on frozen pre-trained experts, its reasoning fidelity can be constrained by their inherent capabilities, occasionally leading to incorrect or hallucinated rationales. Furthermore, the quality of the final prediction is affected by the precision and expressiveness of individual experts. As the ecosystem of multimodal and language models evolves, integrating more advanced experts holds promise for improving both accuracy and robustness without altering the overall framework.
6 Acknowledgment
We thank the reviewers and area chairs for their helpful feedback. This work was supported by the National Institutes of Health (NIH) under other transactions 1OT2OD038045-01, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL211263, and a Capital One Research Award. The views, opinions, and/or findings contained in this article are those of the authors and not of the funding agency.
Appendix A Appendix
A.1 Prompts for Router and Aggregator
Fig. 6 and Fig. 5 illustrate the prompt designs used in our framework, respectively. Fig. 6 presents the expert selection prompt, which identifies skill-specialized experts for the input task context and question. Fig. 5 shows the aggregator prompt, which integrates the outputs from the selected experts and guides LRM to reason over them and generate the final answer.
To generate captions from different experts that explicitly focus on specialized skills, we design and apply skill-specific prompts for each expert. For instance, to obtain detailed image descriptions, we prompt Omnicaptioner-QWen2.5-7B with “You are a helpful assistant focused on providing detailed descriptions and background information for images. Analyze the given image and generate a comprehensive caption that includes the visual style, spatial relationships between elements, texture details, descriptions of the main objects, and relevant world knowledge to enhance understanding." whereas, for concise image summaries, we use the prompt “You are a helpful assistant focused on creating short captions for images. Analyze the provided image and generate a concise caption that highlights the main subject." Similarly, for other experts, we provide clear and direct prompts explicitly highlighting the required skill.
A.2 Evaluation Datasets
We validate our framework across various challenging multimodal tasks, including Video Reasoning (Video-MMMU (Hu et al., 2025b)), Audio QA (MMAU (Sakshi et al., 2024)), 3D Situated Reasoning (SQA3D (Ma et al., 2023b)), and Medical QA (M3D (Bai et al., 2024)). We specifically choose these benchmarks because they represent diverse reasoning complexities, modality interactions, and practical application scenarios.
Video-MMMU (Hu et al., 2025b) provides a robust evaluation of models’ abilities to integrate multimodal educational content and reason across diverse knowledge domains from educational videos. In our evaluation, we test on the full $900$ video reasoning questions.
MMAU (Sakshi et al., 2024) is a benchmark designed to evaluate multimodal audio understanding with curated audio clips paired with human-annotated natural language questions spanning speech, environmental sounds, and music. Evaluation is conducted on a 11K validation set with available ground truth annotations.
SQA3D (Ma et al., 2023b) is a benchmark specifically designed for evaluating situated 3D scene understanding. It presents situations such as "You are standing beside a table", requiring models to reason about 3D spatial relationships between the described viewpoint and surrounding objects. In our evaluation, we use the SQA3D test set, which includes around 33K human-annotated question-answer pairs.
M3D-VQA (Bai et al., 2024) is a benchmark designed for expert-level reasoning over medical data, specifically focusing on 3D CT scans. It consists of natural-language question-answer pairs covering five diagnostic dimensions: plane, phase, organ, abnormality, and location. For evaluation, we sample $500$ question-answer pairs under the closed-ended QA setting.
A.3 Running Efficiency Evaluation
In Tab.˜6, we provide detailed computational cost analysis, including: average expert skills calls and running time (router, captioner and aggregator) as listed in Table A. We observe that Video-MMMU requires, on average, more skills than other tasks. This is because video tasks contain richer multimodal information and therefore need more experts or skills to handle effectively. Additionally, our system allows a flexible trade-off between running time and accuracy by adjusting the size of the expert pool in the router prompt.
Dataset Avg. Skills Call Avg. Running Time
VideoMMMU 2.10 21.10
SQA3D 1.81 18.03
MMAU 1.10 13.31
M3D 1.10 13.81
Table 6: Average skill calls (times) and running time (s) across datasets/benchmarks.
A.4 Comparison With Other Agent Systems
In Tab.˜7, we provide extra results with the recent multimodal multi-agent system. Our MEXA achieves better performance on Video-MMMU and shows more generalizability over different tasks, while other methods are limited to the specific task.
Method Video-MMMU MMAU
VideoMultiAgent Kugo et al. (2025) $62.88$ not applicable
ReAgent-V Zhou et al. (2025) $61.11$ not applicable
MEXA $71.47$ $45.9$
Table 7: Comparison with other Agent-based works.
A.5 Extra Discussion on Expert Overlapping and Conflict
We clarify that our framework addresses redundancy and conflicts in both expert selection and aggregation. Specifically, our MLLM router dynamically selects experts based on their semantic relevance to the given query, significantly reducing unnecessary redundancy from activating all available experts. Additionally, our aggregator leverages contextual reasoning to integrate expert outputs, effectively resolving conflicts by prioritizing information consistency and relevance.
We use two examples to show how our model addresses redundancy and conflicts, respectively.
Redundancy: Consider the query: "Describe the spatial relationship between the table and the chair." Suppose both the 3D Expert and Visual Expert independently provide the response: "The chair is placed next to the table on its left side." In this scenario, our aggregator effectively integrates these overlapping outputs, boosting confidence in the identified spatial relationship ("left") between the table and chair.
Conflicts: Consider the query: "Which object is closer to the door, the chair or the table?" In this scenario, suppose two different experts provide conflicting outputs. The 3D scan expert generates: "The chair is closer to the door.", while the visual expert generates: "The table appears closer to the door." This exemplifies a conflict where two experts offer totally opposing answers. Our aggregator resolves such conflicts by analyzing the query and prioritizes the 3D scan expert, as spatial depth is more accurately captured through 3D geometry than 2D visual appearance.
A.6 Qualitative Examples
Fig 5 and Fig 5 show two qualitative examples from Video-MMMU and SQA3D, respectively. In the Video-MMMU example, our framework effectively selects the most relevant experts, including the video expert and the medical image expert. The aggregator then filters and prioritizes key information extracted by the medical image expert, allowing the reasoning module to accurately fill in the missing information required to answer the question. In the SQA3D example, both the general 3D scene expert and the situated 3D scene expert are activated. Their outputs are jointly considered by the aggregator to produce a coherent answer.
Figure 3: A qualitative example of Video-MMMU.
Figure 4: A qualitative example of SQA3D.
Prompts for Aggregator You are an answerer for a video question answering, audio question answering, 3D situated question answering, or medical visual question answering. Below is information provided by multiple expert modules relevant to solving the question: Expert 1: {Text Descriptions from Expert 1} Expert 2: {Text Descriptions from Expert 2} Expert 3: {Text Descriptions from Expert 3} Using the information above, please select the best answer to the question and provide a brief explanation if needed. Question: "{}"
Figure 5: Prompts for Aggregator.
Prompts for Expert Selection You are an expert multimodal reasoning assistant. Given a multimodal question (e.g., related to video, audio, 3D scenes, medical images, etc.), your task is to select all relevant skills and modalities required to accurately answer the question. Task Type: "{}" Question: "{}" Options: {} Available Skills and Modalities: General Visual Perception A1. Detailed Image Description A2. Medium Image Description A3. Short Image Description Audio Perception B1. Video Subtitle Extraction B2. Audio Description B3. Music Description 3D Visual Understanding C1. 3D Scene Description C2. 3D Situated Context Description Medical Visual Understanding D1. CT Scan Interpretation D2. Medical Image Description OCR/Text Extraction E1. General OCR E2. Poster/Slides Caption E3. PDF Text Extraction Structured Visual Data Interpretation F1. Chart/Plot Description F2. Table Description Mathematics and Geometry Extraction G1. Equation (LaTeX format) G2. Mathematics & Geometry (LaTeX format) Instructions: 1. Only select skill/modality IDs necessary to answer the provided question. 2. Respond strictly with the selected skill IDs, separated by commas. Selected IDs:
Figure 6: Prompts for expert selection.